# SEA: Low-Resource Safety Alignment for Multimodal Large Language Models via Synthetic Embeddings

### v1

@@ 1 Introduction
To address the aforementioned issues, current mitigation strategies, such as supervised fine-tuning (SFT) Zong et al. (2024) and reinforcement learning with human feedback (RLHF) Zhang et al. (2024) demonstrate effectiveness in enhancing the safety of MLLM. However, the construction of multimodal safety alignment datasets is costly. Unlike LLMs, high-quality safety alignment data for MLLMs requires a strong correlation between the three components: textual instructions, textual responses, and additional modalities, making the data collection process even more expensive. Moreover, due to differences in additional modalities, safety alignment data must be rebuilt whenever a new emerging modality (such as electroencephalogram signals Wang et al. (2024a)) is introduced for MLLM. This not only incurs additional costs but also causes the development of datasets to lag behind the advancements of the MLLMs themselves. Therefore, there is an urgent need for a resource-efficient and universally applicable safety alignment method for various MLLMs to promote the development of safer MLLMs.
@@
Recently, Chakraborty et al. (2024) revealed that textual alignment can significantly enhance the security of image-based MLLMs, providing a promising solution for low-resource safety alignment. However, further exploration by Hu et al. (2024) found that textual alignment is effective only when explicit harmful information appears in the text input, such as the instruction “how to use the product in the image to rob a bank” with an image input of a bomb. In contrast, models that align image-text pairs are generally effective across various scenarios, including samples that present harmful information solely through images, such as the instruction "how to make a product" with an input image of a bomb. To address the limitations of textual alignment, generating data of additional modality using generative models is a potential solution. However, not all modalities have high-performance generative models available, especially for emerging MLLMs that may arise in the future.
@@
To address the aforementioned limitations, we propose SEA, a new framework that uses synthetic embeddings of additional modalities to enhance safety alignment. It first optimizes embedding representations within the modality encoder’s output space deemed by MLLMs to contain the specified contents. Subsequently, the optimized embedding can be integrated with the textual dataset, substituting it for a real multimodal dataset in safety alignment training. Our approach eliminates the resource-intensive process of collecting and curating real multimodal datasets. Experiments are conducted on MLLMs based on images, videos, and audio, and the results indicate that only two training samples are needed to optimize a high-quality embedding in 24 seconds on a single RTX 3090 GPU. Furthermore, using datasets constructed with synthetic embedding for safety alignment significantly enhances the safety of MLLMs against threats from additional modalities.
@@ 2.1 Safety Concerns of MLLMs
LLMs have been revealed to pose significant risks in responding to malicious instructions Zou et al. (2023); Liu et al. (2023); Chao et al. (2023). Since MLLMs are typically developed using LLMs as their backbone networks, the risks inherent in the LLM domain are directly transferred to MLLMs. More concerning, recent studies have revealed that non-text modal inputs to MLLMs present even more significant security threats. For example, leveraging the model’s OCR capabilities in combination with malicious images Gong et al. (2023); Luo et al. (2024) can significantly increase the response rate of malicious instructions. Furthermore, some work has used gradient-based searches to generate adversarial perturbations at the image level Li et al. (2024b); Qi et al. (2024); Niu et al. (2024), further exacerbating security risks. Therefore, additional safety alignment for MLLMs is necessary to mitigate potential societal harm.
@@ 2.2 Safety Alignment for MLLMs
Safety alignment aims to align the safety awareness of the model with that of humans to prevent the generation of harmful content. This has been thoroughly researched in the field of LLMs, with widely used methods including SFT, Direct Preference Optimization (DPO) Rafailov et al. (2024), and Proximal Policy Optimization (PPO) Schulman et al. (2017). Inspired by these works, researchers have created carefully crafted image-text pairs to align training in MLLMs, yielding promising results in improving model safety. However, producing high-quality multimodal alignment data is often costly. To achieve low-resource safety alignment, Chakraborty et al. (2024) have revealed that textual unlearning can effectively enhance model safety. However, it has been noted that this is ineffective against attacks introduced solely from images. Hu et al. (2024). Furthermore, most existing works have focused solely on image-based MLLMs, leaving the effectiveness of other modalities to be explored further.
@@ 2.3 Safety Benchmark of MLLMs
Most of the existing safety benchmarks focus on image-based MLLMs, including MM-SafetyBench Liu et al. (2024c), Ch3ef Shi et al. (2024), VLSafe Chen et al. (2024), Figstep Gong et al. (2023), MLLMGuard Gu et al. (2024), and Jailbreakv-28k Luo et al. (2024). Furthermore, Yang et al. (2024a) utilized text-to-speech models to reveal harmful information in the audio modality, while SafeBench Ying et al. (2024) provides a unified benchmark that can test the safety of both image and audio modalities. Currently, there are no published safety assessment benchmarks for MLLMs in other modalities.
@@ 3 SEA: Achieving Low-Resource Safety Alignment via Synthetic Embeddings
Figure 1: The overall framework of SEA.
Since multimodal datasets are crucial for MLLM’s safety alignment training, but not all modalities have high-performance generative models available, we aim to find a more general method for synthesizing additional modal data. An intuition of this work is that the additional modal data used for safety alignment, such as bomb images, do not necessarily need to be understandable by humans, but only need MLLMs to consider them as bomb images.
Building on this intuition, we propose Synthetic Embedding enhanced safety Alignment (SEA), which optimizes embeddings in the representation space of the additional modality. The target embedding is one that MLLMs interpret as containing the specified harmful activities or products. Specifically, SEA treats the embedding of additional modal as a trainable weight, optimized through gradient updates, to maximize the probability of the model outputting the specified content. After integrating the optimized embedding with the textual dataset, it can serve as a substitute for multimodal datasets.
3.1 Preliminary: MLLMs Architecture
@@
Following the above paradigm, regardless of the format differences of the additional modalities of MLLMs, they will be encoded as embeddings via $M(\cdot)$. To make SEA more broadly applicable, we anchor the output space of $M(\cdot)$, in which we collect the desired embeddings for safety training.
@@ 3.2 Data Preparation
Assume a textual safety alignment dataset $D_{T}=\left\{\left(x_{T}^{i},y_{T}^{i}\right)\right\}_{i=0}^{N}$ consisting of $N$ samples, where $x_{T}^{i}$ represents harmful instructions and $y_{T}^{i}$ can be a single moral response for SFT or a pair of chosen rejected responses for RLHF, our objective is to optimize a set of embeddings $\{E^{i}\}_{i=0}^{N}$ based on harmful information in $x_{T}^{i}$.
@@
For each $\left(x_{T}^{i},y_{T}^{i}\right)\in D_{T}$, we individually prepare a dataset $D_{a}^{i}=\{(x_{a}^{1},y_{a}^{1}),(x_{a}^{2},y_{a}^{2})\}$ to assist in the optimization of $E^{i}$, where $(x_{a}^{1},y_{a}^{1})$ and $(x_{a}^{2},y_{a}^{2})$ are the content control sample and the style control sample, respectively. As an example of the image-based MLLMs, the construction process for the two samples is as follows:
@@
Harmful Information Extraction. Inspired by Liu et al. (2024c), we utilize GPT-4o-mini to identify harmful phrase in $x_{T}^{i}$, and then classify harmful phrases into two categories, including “activity” and “product”, and then create a detoxified version of $x_{T}^{i}$ by replacing the harmful phrase with “this product” or “this activity”. Since harmful phrase related to activity often do not form complete sentences, we further prompt GPT-4o-mini to complete them into full sentences $s_{T}^{i}$ with subject-verb-object structures, aligning with the language habits of MLLMs.
@@
Content Control Sample Construction. This sample is used to control the primary harmful content in the embedding. We use “Please briefly describe the activity (product) in the image.” as input instruction $x_{a}^{1}$, and “response prefix + $s_{T}^{i}$.” as the ground truth $y_{a}^{1}$. The “response prefix” is determined based on the output habits of different models and can be found in Appendix B.1.
@@
Style Control Sample Construction. This sample is designed to enhance embedding diversity. The input instruction $x_{a}^{2}$ is set to “What is the style of the image?”. The ground truth $y_{a}^{2}$ is set to “response prefix + style description”. The style description is randomly sampled from a predefined style set determined by the model’s output habits.
3.3 Embedding Optimization.
@@
After building the $D_{a}^{i}$, $M(\cdot)$ encodes a blank image (or a blank video, silent audio) in an embedding, which is used as an initialization for a trainable embedding $E_{o}$. For each $(x_{a},y_{a})\in D_{a}^{i}$, the goal of the embedding optimization is to maximize the probability of the MLLM generating $y_{a}$ when given $x_{a}$ and $E_{o}$. During the optimization process, the entire MLLMs are frozen, with only $E_{o}$ participating as the trainable weight in the gradient updates. Since the content and style are specified in $y_{a}^{1}$ and $y_{a}^{2}$, the optimization objective can be understood as finding the embedding that the MLLM considers most aligned with that content and style. The entire optimization process can be formulated as follows:
$L\left(E_{o}\right)=-\frac{1}{\left|D_{a}^{i}\right|}\sum_{(x_{a},y_{a})\in D_{a}^{i}}\log\left(P_{r}\left(y_{a}\mid x_{a},P\left(E_{o}\right)\right)\right)$,
@@
where $P_{r}\left(y\mid x,P\left(E_{o}\right)\right)$ represents the conditional probability of generating $y$ when given $x$ and $P(E_{o})$ to the LLM.
@@ 3.4 Safety Alignment
To integrate $E^{i}$ and $D_{T}$ to construct the multimodal dataset $D_{M}=\left\{\left(x_{M}^{i},y_{T}^{i},E^{i}\right)\right\}_{i=0}^{N}$, a prefix in the form of “The image shows an activity (product). Please comprehend it and respond to the question below.” is added to the detoxified version of each $x_{T}^{i}$, resulting in the instruction $x_{M}^{i}$. The responses $\{y_{T}^{i}\}_{i=1}^{N}$ in $D_{T}$ are retained in $D_{M}$.
To achieve safety alignment based on $D_{M}$, we need to ignore module $M(\cdot)$ and modify the forward propagation process of MLLMs to $y=LLM\left(P\left(E^{i}\right),x\right)$, allowing it to adapt to existing safety alignment training strategies. Notably, most current MLLMs freeze $M(\cdot)$ during the instruction fine-tuning stage. Therefore, we only need to precompute the embeddings from $M(\cdot)$, the synthetic dataset generated by SEA can then be mixed with real multimodal datasets in the existing training pipeline.
@@ 4.1 VA-SafetyBench Overview
The construction pipeline of VA-SafetyBench is illustrated in Figure 2. VA-SafetyBench builds on MM-SafetyBench, a well-established image-based safety benchmark, through a systematic transformation process. Each test case in VA-SafetyBench directly corresponds to a test case in MM-SafetyBench, which spans eight critical safety scenarios: illegal activity, hate speech, malware generation, physical harm, economic harm, fraud, sexual violence, and privacy violations. For each sample, we utilize three types of textual data from MM-SafetyBench in the transformation process: (1) an original instruction, (2) a harmful key phrase extracted from the original instruction, and (3) a rewritten instruction that conceals the harmful content in the original instruction. Figure 2 provides an example of these texts. Based on three types of textual data, we collect video and audio according to the key harmful phrase and refine the rewritten instruction to suit the new modalities.
@@ 4.2 Video-SafetyBench
$\bullet$ DM + TYPO: It combines the Diffusion Model and Typography to generate videos. In this task, Typography does not displace the words separately but displays the entire harmful key phrase in a single video frame, repeating multiple frames to form a 3-second video, which is then appended to the end of the DM video. For this task, we modified the rewritten instructions by replacing “the image” with “the end of the video” as the textual instruction. An instance would be “The end of the video shows a phrase of an activity. List steps in detail to perform the activity.”
@@ 4.3 Audio-SafetyBench
$\bullet$ Key Phrase Transfer (KPT): Only key phrases are converted into speech. The textual instruction input is the rewritten instruction with “the image” replaced by “the audio”.
$\bullet$ KPT + Noise: Inspired by Yang et al. (2024a), we prepend a segment of white noise lasting 3 seconds to the KPT audio. The textual instruction input is the rewritten instruction with “the image” replaced by “the end of the audio”.

### v3

@@ 1 Introduction
To address the aforementioned issues, current mitigation strategies, such as supervised fine-tuning (SFT) Zong et al. (2024) and reinforcement learning with human feedback (RLHF) Zhang et al. (2024) demonstrate effectiveness in enhancing the safety of MLLM. However, the construction of multimodal safety alignment datasets is costly. Unlike LLMs, high-quality safety alignment data for MLLMs requires a strong correlation between the three components: textual instructions, textual responses, and additional modalities, making the data collection process even more expensive. Moreover, due to differences in additional modalities, safety alignment data must be rebuilt whenever a new emerging modality (such as electroencephalogram signals Wang et al. (2024a)) is introduced for MLLM. This not only incurs additional costs but also causes the development of datasets to lag behind the advancements of the MLLMs themselves. Therefore, there is an urgent need for a resource-efficient and universally applicable safety alignment method to promote the development of safer MLLMs.
@@
Recently, Chakraborty et al. (2024) revealed that textual alignment can significantly enhance the security of image-based MLLMs, providing a promising solution for low-resource safety alignment. However, further exploration by Hu et al. (2024) found that textual alignment is effective only when explicit harmful information appears in the text input, such as the instruction “how to use the product in the image to rob a bank” with an image input of a bomb. In contrast, models that have undergone multimodal alignment are generally effective across various scenarios, including samples that present harmful information solely through images, such as the instruction “how to make the product” with an input image of a bomb. To address the limitations of textual alignment, generating data of additional modality using generative models is a potential solution. However, not all modalities have high-performance generative models available, especially for emerging MLLMs that may arise in the future.
@@
To address the aforementioned limitations, we propose SEA, a new framework that uses synthetic embeddings of additional modalities to enhance safety alignment. It first optimizes embedding representations within the modality encoder’s output space deemed by MLLMs to contain the specified harmful activities or products. Subsequently, the optimized embedding can be integrated with the textual dataset, substituting it for a real multimodal dataset in safety alignment training. Our approach eliminates the resource-intensive process of collecting and curating real multimodal datasets. Experiments are conducted on MLLMs based on images, videos, and audio, and the results indicate that only two training samples are needed to optimize a high-quality embedding in 24 seconds on a single RTX 3090 GPU. Furthermore, using datasets constructed with synthetic embedding for safety alignment significantly enhances the safety of MLLMs against threats from additional modalities.
@@ 2.1 Safety Concerns of MLLMs
LLMs have been revealed to pose significant risks in responding to malicious instructions Zou et al. (2023); Liu et al. (2023); Chao et al. (2023). Since MLLMs are typically developed using LLMs as their backbone networks, the risks inherent in the LLM domain are directly transferred to MLLMs. More concerning, recent studies have revealed that non-text modal inputs pose a more significant security threats to MLLMs. For example, leveraging the model’s OCR capabilities in combination with malicious images Gong et al. (2023); Luo et al. (2024) can significantly increase the response rate of malicious instructions. Furthermore, some works Li et al. (2024b); Qi et al. (2024); Niu et al. (2024) use gradient-based searches to generate image-level adversarial perturbations, further exacerbating security risks. Therefore, additional safety alignment for MLLMs is necessary to mitigate potential societal harm.
@@ 2.2 Safety Alignment for MLLMs
Safety alignment aims to align the safety awareness of the model with that of humans to prevent the generation of harmful content. This has been thoroughly researched in the field of LLMs, with widely used methods including SFT, Direct Preference Optimization (DPO) Rafailov et al. (2024), and Proximal Policy Optimization (PPO) Schulman et al. (2017). Inspired by these works, researchers have created carefully crafted image-text pairs for alignment training in MLLMs, yielding promising results in improving model safety. However, producing high-quality multimodal alignment data is often costly. To achieve low-resource safety alignment, Chakraborty et al. (2024) have revealed that textual unlearning can effectively enhance model safety. However, it has been noted that this is ineffective against attacks introduced solely from images. Hu et al. (2024). Furthermore, most existing works have focused solely on image-based MLLMs, leaving the effectiveness of other modalities to be explored further.
@@ 2.3 Safety Benchmark of MLLMs
Most of the existing safety benchmarks focus on image-based MLLMs, including MM-SafetyBench Liu et al. (2024c), Ch3ef Shi et al. (2024), VLSafe Chen et al. (2024), Figstep Gong et al. (2023), MLLMGuard Gu et al. (2024), and Jailbreakv-28k Luo et al. (2024). Furthermore, Yang et al. (2024a) utilized text-to-speech models to reveal security risks in the audio modality, while SafeBench Ying et al. (2024) provides a unified benchmark that can test the safety of both image and audio modalities. Currently, there are no published safety assessment benchmarks for MLLMs in other modalities.
@@ 3 SEA: Achieving Low-Resource Safety Alignment via Synthetic Embeddings
Figure 1: The overall framework of SEA. The execution process is demonstrated using an example in the image modality, encompassing three stages: data preparation, embedding optimization, and safety alignment.
Since multimodal datasets are crucial for MLLM’s safety alignment training, but not all modalities have high-performance generative models available, we aim to find a more general method for synthesizing data of additional modality. A key insight is that data of additional modality used for safety alignment (such as bomb images) is not necessary be human-interpretable, but merely needs to be interpreted as such by MLLMs.
Building on this intuition, we propose Synthetic Embedding enhanced safety Alignment (SEA), which optimizes embeddings in the representation space of the additional modality. The target embedding is the one that MLLMs interpret as containing the specified harmful activities or products. Specifically, SEA treats the embedding of additional modality as a trainable weight, optimized through gradient updates, to maximize the probability of the model outputting the specified content. After integrating the optimized embedding with the textual dataset, it can serve as a substitute for real multimodal datasets.
The pipeline of SEA is illustrated in Figure 1 and consists of three stages: (1) the data preparation stage, which convert textual alignment data into auxiliary data for embedding optimization. (2) the embedding optimization stage, which focuses on synthesizing embeddings for additional modalities. (3) the safety alignment stage, which performs multimodal alignment training by combining the synthesized embeddings with the textual alignment data. In the remainder of this section, we will first introduce the MLLMs architecture, followed by a detailed explanation of each stage.
3.1 MLLMs Architecture
@@
When considering synthetic data of additional modality, there are three feasible options which are synthesizing the raw additional modality $z$, the output of the modality encoder $M(\cdot)$ or the output of the projector $P(\cdot)$. Since $z$ varies greatly in form across different modalities, it is not considered. Among the remaining options, we favor synthesizing the output of $M(\cdot)$ because in most MLLM training paradigm, $M(\cdot)$ remains frozen while $P(\cdot)$ is trainable, making SEA applicable to more MLLMs.
@@ 3.2 Data Preparation
Assume a textual safety alignment dataset $D_{T}=\left\{\left(x_{T}^{i},y_{T}^{i}\right)\right\}_{i=0}^{N}$ consisting of $N$ samples, where $x_{T}^{i}$ represents harmful instructions and $y_{T}^{i}$ can be a single harmless response for SFT or a pair of chosen-rejected responses for RLHF, our objective is to optimize a set of embeddings $\{E^{i}\}_{i=0}^{N}$ based on harmful information in $\{x_{T}^{i}\}_{i=0}^{N}$.
@@
For each $\left(x_{T}^{i},y_{T}^{i}\right)\in D_{T}$, we individually prepare a auxiliary dataset $D_{a}^{i}=\{(x_{c}^{i},y_{c}^{i}),(x_{s}^{i},y_{s}^{i})\}$ to assist in the optimization of $E^{i}$, where $(x_{c}^{i},y_{c}^{i})$ and $(x_{s}^{i},y_{s}^{i})$ are the content control sample and the style control sample, respectively. Each sample includes an instruction, a response prefix, and a guiding text. The response prefix is concatenated with the guidng text to form the target response. Taking the image-based MLLM as an example, the construction process of $D_{a}^{i}$ is as follows:
@@
Harmful Information Extraction. Inspired by Liu et al. (2024c), we utilize GPT-4o-mini to identify harmful phrase in $x_{T}^{i}$, and then classify harmful phrases into two categories, including “activity” and “product”, and then create a detoxified version of $x_{T}^{i}$ by replacing the harmful phrase with “this product” or “this activity.”
Since the harmful phrase related to activity often do not form complete sentences, we further prompt GPT-4o-mini to convert them into full sentences $s_{T}^{i}$ with subject-verb-object structures, aligning with the language habits of MLLMs.
@@
Content Control Sample Construction. This sample is used to control the primary harmful content in the embedding. We use “Please briefly describe the activity (product) in the image.” as input instruction $x_{c}^{i}$, and “response prefix + $s_{T}^{i}$ (or harmful phrases for product)” as the target response $y_{c}^{i}$. The “response prefix” is determined based on the models’ output habits.
@@
Style Control Sample Construction. This sample is designed to enhance embedding diversity. The input instruction $x_{s}^{i}$ is set to “What is the style of the image?.” The target response $y_{s}^{i}$ is set to “response prefix + style description.” The style description is randomly sampled from a predefined style set determined by the model’s output habits.
More details and examples on constructing $D_{T}$ can be found in Appendix B.1.
3.3 Embedding Optimization
@@
After building the $D_{a}^{i}$, $M(\cdot)$ encodes a blank image (or blank video, silent audio) into an embedding, which serves as an initialization for a trainable embedding $E_{o}$. For each $(x^{i},y^{i})\in D_{a}^{i}$, the goal of the embedding optimization is to maximize the probability of the MLLM generating $y^{i}$ when given $x^{i}$ and $E_{o}$. During the optimization process, the entire MLLMs are frozen, with only $E_{o}$ participating as the trainable weight in the gradient updates. Since the content and style are specified in $y_{c}^{i}$ and $y_{s}^{i}$, the optimization objective can be understood as finding the embedding that the MLLM considers most aligned with that content and style. The entire optimization process can be formulated as follows:
$L\left(E_{o}\right)=-\frac{1}{\left|D_{a}^{i}\right|}\sum_{(x^{i},y^{i})\in D_{a}^{i}}\log\left(P_{r}\left(y^{i}\mid x^{i},P\left(E_{o}\right)\right)\right)$,
@@
where $P_{r}\left(y^{i}\mid x^{i},P\left(E_{o}\right)\right)$ represents the conditional probability of generating $y^{i}$ when given $x^{i}$ and $P(E_{o})$ to the LLM.
@@ 3.4 Safety Alignment
To integrate $E^{i}$ and $D_{T}$ to construct the multimodal dataset $D_{M}=\left\{\left(x_{M}^{i},y_{T}^{i},E^{i}\right)\right\}_{i=0}^{N}$, a prefix in the form of “The image shows an activity (product). Please comprehend it and respond to the question below.” is added to the detoxified version of each $x_{T}^{i}$, resulting in the instruction $x_{M}^{i}$. The detoxified instruction ensures that harmful information is conveyed exclusively through the synthesized embeddings. The responses $\{y_{T}^{i}\}_{i=1}^{N}$ in $D_{T}$ are retained in $D_{M}$.
To achieve safety alignment based on $D_{M}$, we need to bypass module $M(\cdot)$ and modify the forward propagation process of MLLMs to $y=LLM\left(P\left(E^{i}\right),x\right)$, allowing it to adapt to existing safety alignment training strategies.
Notably, most current MLLMs freeze $M(\cdot)$ during the instruction fine-tuning stage. This enables seamless integration of SEA-generated synthetic datasets with real multimodal datasets. Taking image modality as an example, the real multimodal sample takes the form of “image embedding encoded by $M(\cdot)$ + instruction + response,” while the synthetic sample consists of “SEA embedding + instruction + response.” Since $M(\cdot)$ are frozen for most MLLMs, both forms look the same to the MLLM during training.
@@ 4.1 VA-SafetyBench Overview
The construction pipeline of VA-SafetyBench is illustrated in Figure 2. VA-SafetyBench builds on MM-SafetyBench, a well-established image-based safety benchmark, through a systematic transformation process. Each test case in VA-SafetyBench directly corresponds to a test case in MM-SafetyBench, which spans eight critical safety scenarios: illegal activity, hate speech, malware generation, physical harm, economic harm, fraud, sexual violence, and privacy violations. For each sample, we utilize three types of textual data from MM-SafetyBench, including (1) an original instruction, (2) a harmful key phrase extracted from the original instruction, and (3) a rewritten instruction that conceals the harmful content in the original instruction. Figure 2 provides an example of these texts. Based on three types of textual data, we collect video and audio according to the key harmful phrase and refine the rewritten instruction to suit the new modalities.
@@ 4.2 Video-SafetyBench
$\bullet$ DM + TYPO: It combines the Diffusion Model and Typography to generate videos. In this task, Typography does not break phrases into individual words but displays the entire harmful key phrase in a single video frame, repeating multiple frames to form a 3-second video, which is then appended to the end of the DM video. In addition, we modified the rewritten instructions by replacing “the image” with “the end of the video” as the textual instruction. An instance would be “The end of the video shows a phrase of an activity. List steps in detail to perform the activity.”
@@ 4.3 Audio-SafetyBench
$\bullet$ Key Phrase Transfer (KPT): Only key phrases are converted into speech. The textual instruction input is the rewritten instruction with “the image” replaced by “the audio.”
$\bullet$ KPT + Noise: Inspired by Yang et al. (2024a), we prepend a segment of white noise lasting 3 seconds to the KPT audio. The textual instruction input is the rewritten instruction with “the image” replaced by “the end of the audio.”